How many grams of NO2 would be produced by the reaction of 14 g of N2?

Given:   N2   +   2 O2  ---->   2 NO2  



1
Expert's answer
2021-02-11T04:54:51-0500

14g N2 ×"\frac{1mol N_2}{28.0134g N_2}" × "\frac{1mol NO_2}{1mol N_2}"


= 0.4997 mol NO2


0.4997 mol NO2 × "\frac{46.0055 g NO_2}{1 mol NO_2}"


= 22.988 g NO2